On a map, 1 inch equals 13.3 miles. Of two cities are 4.5 inches apart on the map, how far are they acutally apart
Answer by jim_thompson5910(35256)

1/13.3 = 4.5/x

x = 13.3*4.5

x = 59.85


So the two cities are 59.85 miles apart.
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------